# VIN-NBV: A View Introspection Network for Next-Best-View Selection for Resource-Efficient 3D Reconstruction

### v1

@@ 3.3 Design of VIN
The VIN works by reconstructing the scene as a 3D point cloud from the base RGB-D capture $\mathcal{R}_{base}$, alternate techniques like Multiview Stereo can be used in the absence of a depth sensor. We attach simple lightweight features to every point in the point cloud by computing surface normals and the number of views in $I_{base}$ in which each point is visible. Surface normal variance from later down-sampling provides helpful information about the local geometry of the scene. Higher variance can indicate regions of complex geometry that may need additional capture, and lower variance can indicate flatter areas where less capture is sufficient. Point visibility can serve as a confidence measure, where a query view with a high number of points visible in many prior captures is likely less helpful than views where points have a lower visibility.
We create a feature grid for each query view by projecting the featurized 3D point cloud $\mathcal{R}_{base}$ into all $n$ query cameras $C_{q}$. If a pixel on the feature grid does not map back to a point in the reconstructed point cloud, we simply assign a vector with all zeros, otherwise, we use the point features and depth after projection. Per-pixel depth values can help identify depth inconsistencies in the views, which may indicate gaps or holes in the reconstruction. Thus, for every query view, we obtain a feature grid of size 512$\times$512$\times$5. We then downsample these feature grids using a pooling operation to obtain a 256$\times$256 feature map $F_{p}$ for each view. Down-sampling provides an explicit measure of local feature variance and helps reduce overall computation. We concatenate the per-pixel feature variances $F_{v}$ with the feature map $F_{p}$ to provide additional context about local geometric complexity.
@@
We define a convolutional encoder $V_{view}=\mathcal{E}_{\theta}\bigl{(}F_{v}\oplus F_{p}\bigr{)}$ that is applied to the feature grids to transform the local per-pixel information into a global view feature vector $V_{view}$ of size $256\times 1$, $\oplus$ means concatenation. In addition to $V_{view}$, we also compute an empty feature $F_{empty}$ to provide explicit information about reconstruction coverage, allowing the VIN to focus on learning key information on top of this to predict reconstruction improvement. Here, we consider a view pixel to be ”empty” if it does not map back to a point in the reconstruction. To construct $F_{empty}$, we create a hull around the non-empty pixels in our query view and compute the number of empty pixels inside and outside of this hull. Empty pixels inside the hull expose “holes” in the existing reconstruction and can help measure how incomplete the currently captured geometry is. Empty pixels outside the hull correspond to image areas where no part of the current reconstruction is visible. Their count estimates the portion of the view that could reveal entirely new geometry. We concatenate both of these values to create the two-element $F_{empty}$ feature vector. We use $F_{empty}$ as a lightweight proxy for more traditional reconstruction coverage measures.
@@
We also provide the number of base views $F_{base}$ to help indicate at what stage in the capture we are at, since helpful views in the earlier stages look different from helpful views in later stages, and this may change how the model scores different views. We concatenate all of this information with $V_{view}$ and pass it through a MLP $\mathcal{M}(\cdot)$ to predict the final improvement score:
@@
^$\widehat{\mathcal{RRI}}(q)=\mathcal{M}_{\phi}\bigl{(}\mathcal{E}_{\theta}\bigl{(}F_{v}\oplus F_{p}\bigr{)}\oplus F_{empty}\oplus F_{base}).$ (4)
@@ 3.4 Training of VIN
Our objective is to train the VIN to estimate the fitness of a query view, defined as its Relative Reconstruction Improvement ($\mathcal{RRI}(q)$) over the existing base views (as described in Equation 2). To achieve this, we adopt an imitation learning approach. At each acquisition step, we exhaustively compute the $\mathcal{RRI}$ for a set of candidate query views by rendering their RGB-D images and reconstructing the point cloud using the new image along with the previously captured base images. This computed score, which relies on access to the ground-truth 3D model and rendering engine, is referred to as the Oracle RRI. VIN is then trained to predict this Oracle RRI, without access to the rendered image, using only the RGB-D images from previously captured views and the camera parameters of the query view.
@@
However, directly regressing the $\mathcal{RRI}$ proves to be challenging. The model struggles to generalize well across unseen objects and categories. To address this, we reformulate the task as a classification problem by discretizing the $\mathcal{RRI}$ into 15 ordinal classes, where class 0 indicates the least improvement and class 14 the highest. In this formulation, we recognize that misclassifications between distant classes are more detrimental than those between nearby ones. Therefore, even if VIN cannot always predict the optimal next-best view, it should still identify a reasonably good one. To enforce this, we use a ranking-aware classification loss—CORAL [37]. CORAL transforms ordinal labels into a series of binary classification tasks, encouraging predictions that respect the natural order of the labels. This improves model predictions and reduces large misclassification errors.
@@
An additional challenge arises in defining consistent class labels across different stages of acquisition. Early acquisitions often have higher $\mathcal{RRI}$ values because large portions of the scene remain unseen, so view selection has a greater impact. In contrast, later stages typically yield lower $\mathcal{RRI}$ values, as improvements become more incremental, focusing on resolving occlusions and small gaps. Using a fixed class assignment strategy across all stages would misclassify even the best views in later stages into lower-quality categories.
@@
To overcome this, we normalize $\mathcal{RRI}$ values in a stage-independent manner. We group our training data by capture stage, defined by the number of base views, and determine the standard deviation and mean of the $\mathcal{RRI}$ values in these groups. We then take $\mathcal{RRI}$ for each query view and convert it into a z-score based on the number of standard deviations they are from the mean of their group. We soft clip the z-scores using a tanh function, to prevent extreme outliers, and then group views into 15 dynamically sized bins, ensuring a similar number of samples in each bin. We use these bins as our final class labels.
3.5 Implementation Details
Model Implementation. We implement our model using Pytorch [38] and conduct training using Pytorch Lightning [39]. We leverage Pytorch3D[40] to help create, and render our point clouds. Our convolutional encoder has 4 layers and hidden dimension size of 256. Our rank MLP has 3 layers with a hidden dimension size of 256 and uses a CORAL [37] layer as its final layer so that we can use the CORAL loss during training. We use the open-source implementation of this loss and other necessary components provided by the authors on GitHub.
Point Cloud Projection. To project our point clouds to query views, we use the Pytorch3D[40] PointsRasterizer with a points per pixel setting of 1 and a fixed radius of 0.01. Since the rasterizer uses a normalized coordinate system and our voxel down sampling leads to evenly spaced points we find that this fixed radius size is sufficient. During the projection process for the point cloud we save a mapping from pixel to point allowing us to determine per pixel feature vectors to construct a feature grid. We do all projection in batch.
Model Training. We train our model for 60 epochs using the AdamW[41] optimizer and a cosine annealing[42] learning rate scheduler. For our learning rate we used 1e-3 and train for approximately 24 hours on four A6000 GPUs. During training time we treat each set of next views for one specific object as a single batch. Batching by object means we only need to reconstruct one point cloud for an entire batch and can project it to all candidate views in one go, saving memory and computation.
Evaluation Data Preparation. The authors of GenNBV [25] resized the house objects from OmniObject3D [43] to better fit their problem context. The resized objects were made available by the authors on the project GitHub and we used them to calculate the exact scale factor applied to the original OmniObject3D [43] houses. We use this scale factor to bring our chamfer distance values computed on the unscaled objects into the correct scale for direct comparison.
@@ 4 Experiments
We evaluate the effectiveness of VIN-NBV and compare it to existing NBV approaches to select viewpoints enabling high-quality 3D reconstructions on standard datasets. We specifically focus on scenarios where a robot can acquire only a few images or its motion time is limited.
Datasets. We follow the same train-test protocol from GenNBV[25] by training on a modified subset of Houses3K [18], and testing on the house category from OmniObject3D [43]. We also evaluate our method on additional classes (see Section 4.3). We begin the acquisition process by selecting two base views from 120 renders of the object. The first base view is randomly selected, and the second one is the view with the closest camera position to the first.
@@
NBV Policy Accuracy (cm) $\downarrow$
@@
Baselines. We compare our VIN-NBV policy against several existing 3D-based next-best-view (NBV) approaches, including ScanRL [18], GenNBV [25], and ActiveRMap [44]. Since the evaluation code and model weights of GenNBV[25] are unavailable, we did our best to reproduce their evaluation framework to directly compare with their reported results in the paper. We also perform a direct visual comparison to the reconstruction results provided in their paper, as seen in Fig. 3. Additionally, we include two baseline methods: Coverage NBV and Oracle NBV.
Coverage NBV shares the same greedy view selection strategy as VIN-NBV, but instead of using VIN’s predicted $\mathcal{RRI}$ score (line 9 in Algorithm 3.2), it relies on a coverage-based fitness function, $Cov(q)=W\times H-\sum_{u=1}^{H}\sum_{v=1}^{W}\mathbbm{1}(C_{q}(\mathcal{R}_{base})_{u,v})$, where $W\times H$ is the image resolution and $\mathbbm{1}(\cdot)$ is an indicator function. This function estimates the number of empty pixels in the query view $C_{q}$ when rendering the current reconstructed 3D point cloud $\mathcal{R}_{base}$. A higher $Cov(q)$ score suggests that the query view covers more previously unseen areas, making it a potentially valuable viewpoint.
Similarly, the Oracle NBV baseline is a variant of the VIN-NBV policy, where the fitness function (line 9 in Algorithm 3.2) is replaced with the ground-truth Relative Reconstruction Improvement ($\mathcal{RRI}$) of each query view. This method assumes access to the complete ground-truth reconstruction, enabling direct computation of $\mathcal{RRI}$ as defined in Equation 2.
@@
Metrics. Since our goal is to improve reconstruction quality, we use chamfer distance as our main accuracy metric, as it can capture fine-grained information at the point level. For each object, we calculate the chamfer distance between the reconstructed and the ground truth point clouds. We report the average accuracy (chamfer distance) in centimeters across all objects in Omniobject3D houses [43].
(a) Number of Acquisitions
(b) Time Traversed
@@
Figure 4: Quantitative evaluation of NBV policies on OmniObject3D [43], showing average reconstruction error (in cm) under constraints on (a) number of acquisitions and (b) time traversed.
4.1 Evaluation with Limited Number of Acquisitions
@@
To enable direct comparison with GenNBV and ScanRL [45, 18], we evaluate our policy on the OmniObject3D [43] houses dataset, limited to 20 captures. VIN-NBV outperforms state-of-the-art NBV policies (Table 1), achieving an average reconstruction error of 0.20 cm on houses, compared to 0.33 cm for GenNBV [25] and 0.37 cm for ScanRL. A visual comparison is presented in Fig. 3, modified directly from the visuals presented in the GenNBV [25] paper, shows VIN-NBV retains finer detail, whereas GenNBV and ScanRL produce blurrier results despite good coverage.
@@
In Figure. 4(a) we compare the reconstruction accuracy during the intermediate stages of acquisition between VIN-NBV with the coverage baseline ‘Cov-NBV’ and the ‘Oracle NBV’. We observe that large improvements with respect to the coverage baseline are observed during the initial stages of capture, while towards the later stages, both policies result in similar reconstruction quality. Since the model weights and the evaluation code of the GenNBV [25] paper are unavailable, we were unable to evaluate this method during the intermediate stages of the capture process.
@@
Ablation: Importance of coverage feature $F_{empty}$ in VIN. To understand the importance of the coverage feature we provide with $F_{empty}$, we run an ablation that evaluates the VIN-NBV policy using a VIN model trained without this information. In Fig. 4(a), we see that the absence of coverage features makes no difference during earlier acquisition stages but makes a substantial difference during the later stages. This explains why VIN-NBV achieves large gains over Cov-NBV early on, converging to a similar solution during later stages where coverage is helpful. Although VIN focuses on predicting reconstruction quality, the coverage feature is still a helpful indicator, which is otherwise hard to learn implicitly during training.
4.2 Evaluation with Time-Limited Robot Motion
@@
In our time-limited setting, we evaluate our policy under varying constraints on time in motion to understand its efficiency under strict budgets, simulating many critical time-sensitive applications. We assume that the robot is a drone equipped with a depth sensor and camera and travels at a constant velocity of 4 mph, a typical speed for consumer-grade drones. We assume that the drone always takes a straight-line path to the next capture position and determine how far it can travel under different time limits. We use the same OmniObject3D [43] evaluation data as before and scaled all objects to a size of 14 meters to simulate a more realistic capture setting.
@@
Adapting NBV policies for time-limited acquisitions. At each step during capture, our robot checks how far it can still travel and finds all potential next views that are within range. If our robot is not able to find a next view within range, it stops capture and computes the chamfer distance of the reconstruction with the current captures. We use our VIN-NBV policy to select the next views and do not impose a limit on the number of images that the robot can capture in this setting.
@@
We show our results from the time-limited setting in Fig.  4(b). Under all time constraints, our method outperforms the coverage baseline, with the largest improvement happening under the strictest time constraints. During the 15 second time limit our method outperforms the coverage baseline by $\sim$25%. In Figure. 5, we visualize the reconstructed 3D scene using our VIN-NBV policy and the coverage baseline (Cov-NBV) for varying time constraints, illustrating that with more available time, the policy focuses on capturing views that can fill missing regions and holes in the reconstruction.

### v3

@@ 3.3 Design of VIN
The View Introspection Network (VIN) estimates the utility of a candidate viewpoint by predicting its Relative Reconstruction Improvement (RRI) given the acquisition history and current reconstruction. There are three stages: first it reconstructs the scene, then it featurizes its geometry, and finally it encodes features to predict improvement.
From the current set of base RGB-D images $I_{base}$ and camera parameters $C_{base}$, we reconstruct a 3D point cloud $\mathcal{R}_{base}$. If a depth sensor is unavailable, monocular depth estimation or Multi-View Stereo can be substituted. Each point in $\mathcal{R}_{base}$ is enriched with surface normals, visibility count, and depth values. Surface normals estimate local geometry, with variance across neighboring points indicating complex surfaces that may require additional captures; low variance indicates planar regions that need less captures. Visibility count tracks the number of views from $I_{base}$ in which a point is observed. Points frequently seen in many views are less informative for new captures; points rarely seen may be more informative. Depth values provide distance information for projected pixels, helping to detect surface inconsistencies and missing regions.
For each candidate view $C_{q}$ the enriched point cloud is projected into its image plane, forming a feature grid (size $\times512$$\times$5). Here each pixel contains surface normals, visibility count, and depth after projection. We downsample this grid to $\times256$ via pooling to reduce computation and compute per-pixel variance to capture local geometric complexity. The down sampled feature grid is represented as $F_{p}$ and the per-pixel variances are $F_{v}$.
@@
We also compute an empty feature $F_{empty}$ to provide explicit information about reconstruction coverage. This allows the VIN to focus on learning key information on top of coverage to predict reconstruction improvement. We consider a pixel to be ”empty” if no point in $\mathcal{R}_{base}$ projects to it. We form a hull around non-empty pixels and count empty pixels inside and outside of the hull. Empty pixels inside the hull expose “holes” in the existing reconstruction and empty pixels outside the hull indicate potentially unseen geometry. We concatenate both of these values to create the two-element $F_{empty}$ feature vector. We use $F_{empty}$ as a lightweight proxy for more traditional reconstruction coverage measures.
@@
To encode the features we define a convolutional encoder $V_{view}=\mathcal{E}_{\theta}\bigl{(}F_{v}\oplus F_{p}\bigr{)}$ that is applied to the feature grids and variances to transform the local per-pixel information into a global view feature vector $V_{view}$ of size $256\times 1$, where $\oplus$ means concatenation. In addition to $V_{view}$, we also provide the number of base views $F_{base}$ to help indicate at what stage of the capture we are in; since helpful views in the earlier stages look different from helpful views in later stages, this may change how the model scores different views. We concatenate all of this information with $V_{view}$ and pass it through a MLP $\mathcal{M}(\cdot)$ to predict the final improvement score:
@@
$\widehat{\mathcal{RRI}}(q)=\mathcal{M}_{\phi}\bigl{(}\mathcal{E}_{\theta}\bigl{(}F_{v}\oplus F_{p}\bigr{)}\oplus F_{empty}\oplus F_{base}).$ (4)
@@ 3.4 Training of VIN
Our objective is to train the VIN to estimate the fitness of a query view, defined as its Relative Reconstruction Improvement ($\mathcal{RRI}(q)$) over the existing base views (as described in Equation 2). To achieve this, we adopt an imitation learning approach. For each object we sample and render a set of 120 RGB-D views that surround it, then at each step we exhaustively compute the $\mathcal{RRI}$ for every query view. At each step we reconstruct the point cloud using the new image along with the previously captured base images. This computed score, which relies on access to the ground-truth 3D model and rendering engine, is referred to as the Oracle RRI. VIN is then trained to predict this Oracle RRI, without access to the rendered image, using only the RGB-D images from previously captured views and the camera parameters of the query view.
@@
Directly regressing the $\mathcal{RRI}$ proves to be challenging. The model struggles to generalize across unseen objects and categories. To address this, we reformulate the task as a classification problem by discretizing the $\mathcal{RRI}$ into 15 ordinal classes, where class 0 indicates the least improvement and class 14 the highest. In this formulation, we recognize that misclassifications between distant classes are more detrimental than those between nearby ones. Therefore, even if VIN cannot always predict the optimal next-best view, it should still identify a reasonably good one. To enforce this, we use a ranking-aware classification loss: CORAL [4]. This loss transforms ordinal labels into a series of binary classification tasks, encouraging predictions that respect the natural order of the labels. This improves the model predictions and reduces large misclassifications.
@@
An additional challenge arises in defining consistent class labels across different stages of acquisition. Early acquisitions often have higher $\mathcal{RRI}$ values because large portions of the scene remain unseen, so view selection has a greater impact. In contrast, later stages typically yield lower $\mathcal{RRI}$ values, as improvements become more incremental, focusing on resolving occlusions and small gaps. Using a fixed class assignment strategy across all stages would misclassify even the best views in later stages as lower-quality.
@@
To overcome this, we normalize $\mathcal{RRI}$ values in a stage-independent manner. We group our training data by capture stage, defined by the number of base views, and determine the standard deviation and mean of the $\mathcal{RRI}$ values in these groups. We then take $\mathcal{RRI}$ for each query view and convert it into a z-score based on the number of standard deviations they are from the mean of their group. We soft-clip the z-scores using a tanh function, to prevent extreme outliers, and then group views into 15 dynamically sized bins, ensuring a similar number of samples in each bin. We use these bins as our final class labels.
@@ 4 Experiments
NBV Policy Chamfer Distance (cm) $\downarrow$
@@
We evaluate and compare VIN-NBV to existing NBV approaches that select viewpoints for high-quality 3D reconstructions on standard datasets. We focus on scenarios where an agent can acquire only a few images or has a limited motion time.
Datasets. We follow the same train-test protocol from GenNBV[5] by training on a modified subset of Houses3K [29], and testing on the house category from OmniObject3D [41]. We also evaluate on the dinosaur, truck, and animal classes from OmniObject3D [41] to show the generalization beyond the training data (see Section 4.4). Each object is rendered from 120 views and we begin the acquisition process by selecting the first base view randomly and then set the second base view to be the one closest to the first.
Baselines. Prior NBV approaches explicitly ”look ahead” from hypothetical camera positions by evaluating visibility of the current scene model, either volumetrically by ray-casting through an occupancy/TSDF grid to estimate information gain [19, 15], or via rendering-style approximations of coverage from a discrete set of viewpoints [6], with related ray/visibility accounting in sensor-aware formulations [30] and quality-driven surface methods [40]. Following this approach, we design the Coverage-NBV (Cov-NBV) baseline that scores a candidate view by projecting the current 3D reconstruction into its image plane.
Since our primary goal is to show that our proposed Relative Reconstruction Improvement (RRI) criterion, predicted by the View Introspection Network (VIN), is better than the coverage-based criterion, we use the same sampling-based sequential greedy best next view selection strategy for both VIN-NBV and Cov-NBV. We simply replace the $\mathcal{RRI}$ score (line 9 predicted by VIN in Algo. 1), with a coverage-based fitness function eq. 5 where $W\times H$ is the image resolution and $\mathbbm{1}(\cdot)$ is an indicator function.
$Cov(q)=W\times H-\sum_{u=1}^{H}\sum_{v=1}^{W}\mathbbm{1}(C_{q}(\mathcal{R}_{base})_{u,v})$ (5)
This function estimates the number of empty pixels in the query view $C_{q}$ when rendering the current reconstructed 3D point cloud $\mathcal{R}_{base}$. A higher $Cov(q)$ score suggests that the query view covers more previously unseen areas, making it a potentially valuable viewpoint. This pixel-space surrogate correlates with reconstruction completeness and serves as a strong, transparent baseline. It is sensor-agnostic, cheap to compute with standard z-buffer rendering, and captures the same “look-ahead visibility” principle that underlies volumetric information-gain methods—without requiring full voxel ray marching.
We also compare our VIN-NBV policy against several existing 3D-based next-best-view (NBV) approaches, including ScanRL [29], GenNBV [5], and ActiveRMAP [44]. Since the model weights of GenNBV[5] are unavailable, we compare with their reported results in the paper directly. We also perform a direct visual comparison to the reconstruction results provided in their paper, as seen in Fig. 3.
To understand the upper bound performance of our methods in the optimal case we propose Oracle NBV: a variant of the VIN-NBV policy that selects the best image at each step where the fitness function (line 9 in Algorithm 1) is replaced with the ground-truth Relative Reconstruction Improvement ($\mathcal{RRI}$) of each query view. This method assumes access to the complete ground-truth reconstruction, enabling direct computation of $\mathcal{RRI}$ as defined in Eq. 2.
@@
Metrics. Since our goal is to improve reconstruction quality, we use the Chamfer Distance as our main accuracy metric, as it can capture fine-grained information at the point level. We also include the coverage percentage and F1 score curves to show the change in reconstruction completeness and quality as more acquisitions are made. For each object, we calculate the metrics between the reconstructed and the ground truth point clouds. We report the average accuracy (Chamfer Distance) in centimeters across all objects in Omniobject3D houses [41].
@@
Figure 4: Evaluation with Limited Acquisitions. Average reconstruction error (in cm) on OmniObject3D [41] houses under constraint on number of acquisitions. VIN-NBV consistently outperforms coverage-based NBV policies.
4.1 Evaluation with Limited Acquisitions
@@
To enable direct comparison with GenNBV and ScanRL [32, 29], we evaluate our policy on the OmniObject3D [41] houses dataset, limited to 20 captures. VIN-NBV outperforms state-of-the-art NBV policies (Table 1), achieving an average reconstruction error of 0.20 cm on houses, compared to 0.33 cm for GenNBV [5] and 0.37 cm for ScanRL. A visual comparison is presented in Fig. 3, modified directly from the final reconstructions presented in the GenNBV [5] paper, and shows VIN-NBV retains fine detail and overall good reconstruction quality whereas GenNBV and ScanRL produce blurrier results despite good coverage, likely due to the authors’ use of mesh renderings as opposed to dense clouds.
@@
In Fig. 4 we compare the reconstruction accuracy during the intermediate stages of acquisition between VIN-NBV with the coverage baseline ’Cov-NBV’. We observe that large improvements with respect to the coverage baseline happen during the initial stages of capture, while towards the later stages, both policies result in similar reconstruction quality. Since the model weights and the evaluation code of the GenNBV [5] paper are unavailable, we were unable to evaluate this method during the intermediate stages of the capture process. The gap between the ’Oracle NBV’ and our method shows that there is still room for improvement, especially in the early stages.
(a) Coverage % over captures
(b) F1 score over captures
Figure 5: 5(a) mean coverage percentage and 5(b) F1 score across all OmniObject3D house objects at each acquisition step.
@@
Ablation: Coverage feature $F_{empty}$ in VIN. To understand the importance of the coverage information we provide with $F_{empty}$, we run an ablation that evaluates the VIN-NBV policy using a VIN trained without $F_{empty}$. In Fig. 4, we see that the absence of coverage features makes little difference during the earlier acquisition stages, but is substantial in the later ones. This explains why VIN-NBV achieves large gains over Cov-NBV early on, converging to a similar solution during later stages where coverage is helpful. Although VIN focuses on predicting reconstruction quality, the coverage feature is still a helpful indicator, which is otherwise hard to learn implicitly during training.
4.2 Evaluation with Time-Limited Agent Motion
@@
In our time-limited setting, we evaluate our policy under varying constraints on time in motion to understand its efficiency under strict budgets, simulating time-sensitive applications. We assume that the agent is a drone equipped with a depth sensor and camera and travels at a constant velocity of 4 mph, a typical speed for consumer-grade drones. We assume that the drone always takes a straight-line path to the next capture position and determine how far it can travel under different time limits. We use the same OmniObject3D [41] evaluation data as before and scaled all objects to a size of 14 meters to simulate a more realistic capture setting.
@@
Adapting NBV policies for time-limited acquisitions. At each step during capture, our agent checks how far it can still travel and finds all potential next views that are in range. If our agent is not able to find a next view in range, it stops capturing and computes the Chamfer Distance of the reconstruction with the current captures. We use our VIN-NBV policy to select the NBVs and don’t impose a limit on the number of images that the agent can capture.
Figure 6: Evaluation with Time-Limited Agent Motion. Average reconstruction error (in cm) on OmniObject3D [41] houses under constraint on time traversed. VIN-NBV consistently outperforms coverage based NBV policies.
@@
We show results from the time-limited setting in Fig.  6. Under all time constraints, our method outperforms the coverage baseline, with the largest improvement happening under the strictest time constraints. During the 15-second limit, our method beats the coverage baseline Chamfer Distance by $\sim$25%. In Fig. 9, we visualize the reconstructed 3D scene using our VIN-NBV policy and the coverage baseline (Cov-NBV) for varying time constraints, illustrating that with more time, the policy focuses on capturing views that can fill missing regions and holes in the reconstruction.
Figure 7: Evaluation with Collision Avoidance. Comparison of average reconstruction error (in cm) on OmniObject3D [41] houses with collisions avoidance planning. VIN-NBV performance is similar while Cov-NBV performance decreases.
4.3 Collision Avoidance
Currently our capture agent assumes that it can perform straight line travel between the current position and the selected next view. This approach does not take into account collisions that may occur from following these straight line trajectories. We demonstrate the ability of our method to handle custom collision criteria by running the evaluation in a setting that requires the selected NBV to have a straight line path that is at least 1 meter away from the reconstruction at all times. In Fig. 7, we see that even with this additional collision criterion integrated into our NBV-policy, it still performs well and outperforms the Cov-NBV baseline.
(a) Dinosaurs
(b) Toy Motorcycles
(c) Toy Animals
Figure 8: Generalization Across Object Categories. We graph the average Chamfer Distance at each capture stage in three additional object classes namely dinosaurs, toy motorcycles, and toy animals. We compare VIN-NBV to a coverage baseline for 20 captures.